OpenStudy (anonymous):

square binomial of x2 + 18x + 81

OpenStudy (anonymous):

are you asked to factor this as well?

OpenStudy (anonymous):

note that, its in the form: a^2+2ab+b^2, which can be written as (a+b)^2.

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

example: x^2+6x+9, a=x, b = 3. so, x^2+2(3)(x)+(3)^3=(x+3)^2.

OpenStudy (theviper):

\[\Huge{\color{red}{x^2+18x+81}}\]\[\implies x^2+9x+9x+81\]\[\implies x(x+9)+9(x+9)=??\]

OpenStudy (anonymous):

or similarly: (x+___)(x+____) put in the blanks, pair factor of 81, which when you add you get 18.

OpenStudy (theviper):

Understand @janirish ??

OpenStudy (anonymous):

yes i understand @TheViper

OpenStudy (theviper):

yw:)